Draw the structure of 4-fluorocycloheptanol.
Predict the product of the reaction below: H,SO + H2O H3C-CH2-CH2 heat
Predict the product of the reaction below: OH H,SO H3C-CH2-CH-CH2-CH3 + H2O heat
Answer #1

HO 4-fluorocycloheptanol H OH H2S04H .C=CH2 + H2O heat OH H2CH2C-C-CH2CH3 42804 HyC-CH2 CH=CH-CH3 + H2O heat